OpenStudy (anonymous):

6. Starting from 1.5 miles away, a car drives toward a speed checkpoint and then passes it. The car travels at a constant rate of 53 miles per hour. The distance of the car from the checkpoint is given by d = |1.5 – 53t|. At what times is the car 0.1 miles from the checkpoint? Calculate your answer in seconds. (1 point) 95.1s and 108.7s 10.2s and 101.9s 108.7s and 10.2s 95.1s and 10.2s

OpenStudy (anonymous):

d=|1.5−53t| where d is in miles and t is in hours. first divide both sides by 53 to get: d53=|1.553−t| this can be split into two portions: (1) for t <= 1.5/53 hours we can write: d53=1.553−t t=1.5−d53 (2) for t > 1.5/53 hours we can write: d53=−(1.553−t)=t−1.553 t=1.5+d53 now just stick the value d=0.1 into (1) and (2) above to get your answer (remember to multiply the answers you get by 3600 to convert hours to seconds).

OpenStudy (anonymous):

I'm screwed for the test!

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

this one of the questions I have to do on the test. The teacher gives us a couple of questions from the test to show us what we need to know.